OpenStudy (anonymous):

A volcano shaped like a cone has a diameter of 10 km and height of 4 km. Sam estimates that 1/4 of the volcano js filled with lava. About how much lava is inside the volcano? Use 3 for pi.

OpenStudy (anonymous):

Volume of a cone = (1/3)*pi*r^2*h r=d/2 so solve for the volume and then divide your answer by four because only a fourth of the volcano is filled with lava does that explain it?

OpenStudy (anonymous):

so 26 km ^3?

OpenStudy (anonymous):

if you are using pi=3 then its something like this (1/3)*3*(10/2)^2*4=100km^3 now divide that by 4 and you get 25km^3

OpenStudy (anonymous):

i used a calc and got 26

OpenStudy (anonymous):

but mk.

OpenStudy (anonymous):

because the pi doesn't equal 3, but when we use 3 for pi its how i showed you above using pi=3.14 the answer is 26.1799

OpenStudy (anonymous):

Thanks:)

OpenStudy (anonymous):

np